Consider the reaction

8H2S(g)+4O2(g)→8H2O(g)+S8(g)8H2S(g)+4O2(g)→8H2O(g)+S8(g)

Δ[H2S]/ΔtΔ[H2S]/Δt = -0.022 M/sM/s

You may want to reference (Pages 587 - 592) Section 14.3 while completing this problem.

Part A

Find Δ[O2]/ΔtΔ[O2]/Δt.

Express your answer to two significant figures and include the appropriate units.

Part B

Find Δ[H2O]/ΔtΔ[H2O]/Δt.

Express your answer to two significant figures and include the appropriate units.

Part C

Find Δ[S8]/ΔtΔ[S8]/Δt.

Express your answer to two significant figures and include the appropriate units.

Part D

Find the rate of the reaction.

Express your answer to two significant figures and include the appropriate units.
Answer #1

rate of reaction is given as:
rate = sign * (1/x) * Δ[concentration]/Δt

sign is negative for reactant and positive for product
x is stoichiometric coefficient in chemical equation

In terms of H2S:
rate = (-1/8) * Δ[H2S]/Δt
rate = (-1/8) * (-0.022 M/s)
= 2.75*10^-3 M/s


a)
In terms of O2,
Rate = (-1/4) * Δ[O2]/Δt
2.75*10^-3 M/s = (-1/4)*Δ[O2]/Δt
Δ[O2]/Δt = 0.011 M/s
Answer: 0.011 M/s

b)
In terms of H2O:
rate = (1/8) * Δ[H2O]/Δt
2.75*10^-3 M/s = (1/8) * Δ[H2O]/Δt
Δ[H2O]/Δt = 0.022 M/s
Answer: 0.022 M/s

c)
In terms of S8:
rate = Δ[S8]/Δt
2.75*10^-3 M/s = Δ[S8]/Δt
Δ[S8]/Δt = 2.75*10^-3 M/s
Answer: 0.0028 M/s

D)
In terms of H2S:
rate = (-1/8) * Δ[H2S]/Δt
rate = (-1/8) * (-0.022 M/s)
= 2.75*10^-3 M/s
Answer: 0.0028 M/s